# Learning to Route: A Rule-Driven Agent Framework for Hybrid-Source Retrieval-Augmented Generation

### v1

@@ 5.6. Analysis of Path Utilization
This experiment examines how routing affects the effectiveness of individual augmentation paths. In Figure LABEL:fig:PDTLABEL:fig:PDW, the bar plots compare two settings: the accuracy obtained when all queries are forced through a single augmentation path (orange), versus the accuracy of the same path when it is selectively chosen by our routing mechanism (blue). In other words, the comparison isolates the effect of routing on path quality, independent of the overall distribution of queries across paths. We observe that across all datasets, the selectively chosen paths consistently achieve higher accuracy than their single-path baselines. For example, on TATQA, the hybrid path shows a clear improvement when only the most suitable queries are routed to it, while the DB and Doc paths also see accuracy gains under selective use. On FinQA and WikiQA, similar trends hold: the DB, Doc, and Hybrid paths all yield higher accuracy under routing than when used in isolation. These results demonstrate that routing not only improves the overall system accuracy but also enhances the effectiveness of each individual augmentation strategy by matching queries with the path best aligned to their information needs. In particular, this highlights that the proposed mechanism avoids inefficient “one-size-fits-all” usage of any single augmentation path, instead leveraging complementarity in a targeted and adaptive manner.
@@ 5.7. Investigation of Path Selection
5.8. Case Study
@@
6. Conclusion
Appendix A Rule Updates

### v2

@@ 5.6. Analysis of Path Utilization
This experiment examines how routing affects the effectiveness of individual augmentation paths. In Figure LABEL:fig:PDTLABEL:fig:PDW, the bar plots compare two settings: the accuracy obtained when all queries are forced through a single augmentation path (orange), versus the accuracy of the same path when it is selectively chosen by our routing mechanism (blue). This comparison isolates the effect of routing on path quality, independent of the overall distribution of queries across paths. We observe that across all datasets, selectively chosen paths consistently achieve higher accuracy than their single-path baselines. For instance, on TATQA as well as FinQA and WikiQA, the DB, Doc, and Hybrid paths all show clear accuracy gains when routing assigns only the most suitable queries to them. These results demonstrate that routing not only improves overall system accuracy but also enhances each individual augmentation strategy by avoiding inefficient one-size-fits-all usage and instead leveraging path complementarity in a targeted and adaptive manner.
@@ 5.7. Investigation of Path Selection
6. Conclusion
Appendix A Case Study
@@
Figure 7. Initial Rule
Figure 8. TATQA Updated Rule
Figure 9. FINQA Updated Rule
Figure 10. WIKIQA Updated Rule
Figure 11. Prompt for Rule-Driven Routing Agent
Figure 12. Prompt for Rule-Driven Routing Agent
Appendix B Rule Updates